x + y + z = 200
3x + 5y + 4z =870
y=4z
You can use substitution solve the system above. Substitute 4z for y in the first two equations. What new system in two variables do you get?

Answer by TimothyLamb(4379)

---
x + y + z = 200
3x + 5y + 4z = 870
y = 4z
---
put the system of linear equations into standard form
---
x + y + z = 200
3x + 5y + 4z = 870
y - 4z = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 50
y= 120
z= 30
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php